OpenStudy (anonymous):

An advertisement states that a television originally priced at three hundred dollars has now been reduced by forty percent. What is the new price

Parth (parthkohli):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Best Answer' button on the right!

Parth (parthkohli):

I am OwlRed. Nice to meet you!

Parth (parthkohli):

So the reduction price would be \(300 - {(0.40 \times 300)}\)

Parth (parthkohli):

I mean the price after reduction ^

OpenStudy (anonymous):

So what is the answer?

Parth (parthkohli):

can you calculate it by using a calculator?

OpenStudy (anonymous):

what do i have to do

OpenStudy (anonymous):

compute \(300\times .6\) and see what you get

OpenStudy (anonymous):

I got 180

OpenStudy (anonymous):

because if it is 40% off you are paying 60% of the price and 60% of 300 means \[300\times.6\]

OpenStudy (anonymous):

yup

Parth (parthkohli):

Oh anyway I have to admit that it was an OwlFred fail :|